# SteerPose: Simultaneous Extrinsic Camera Calibration and Matching from Articulation

### v1

@@ 3.2 Joint Extrinsic Camera Calibration and Matching
The geometric consistency loss $\mathcal{L}_{\mathrm{geom}}$ evaluates whether the combination of the current $R$ and the estimated matches yields a valid solution of the relative translation, after enforcing collinearity and coplanarity constraints  [35, 19]. For all $N$ corresponding 2D pose pairs under current rotation $R$, we transform their positions into a normalized coordinate system. By stacking linear equations derived from collinearity and coplanarity constraints [19], we can obtain a linear system of the form $A\begin{bmatrix}x_{1}&\dots&x_{N}&\bm{t}\end{bmatrix}^{\top}=\bm{0}$, where unknowns $x_{i}$ are 3D poses recovered from the collinearity constraint and $\bm{t}$ is determined up to scale and sign ambiguity. This sign ambiguity is resolved via a chirality check [35].
@@ 5.1 Experimental Setups
We assess the camera pose error using rotation error $E_{R}$ and translation error $E_{\mathbf{t}}$, both expressed as angular differences in degrees, and report the median projection error $E_{\mathrm{2D}}$ in pixels between the reprojected 3D keypoints and their corresponding 2D observations. For the matching evaluation, we measure the precision $P$ using the ground-truth instance correspondences across views. We further evaluate the precision of the estimated camera pose using three metrics: the relative rotation accuracy (RRA), relative translation accuracy (RTA), and area under curve (AUC) [31]. For a given threshold $\tau$, AUC@$\tau$ is computed as the minimum of RRA@$\tau$ and RTA@$\tau$.
@@
RRA@30$\uparrow$ RTA@30$\uparrow$ AUC@30$\uparrow$ RRA@30$\uparrow$ RTA@30$\uparrow$ AUC@30$\uparrow$ RRA@30$\uparrow$ RTA@30$\uparrow$ AUC@30$\uparrow$
@@
RRA@20$\uparrow$ RTA@20$\uparrow$ AUC@20$\uparrow$ RRA@20$\uparrow$ RTA@20$\uparrow$ AUC@20$\uparrow$ RRA@20$\uparrow$ RTA@20$\uparrow$ AUC@20$\uparrow$
@@ 5.2 Comparison with Existing Methods
Figure 4: Visualization for multi-view calibration results. The ground-truth cameras are in black, the initial camera poses after registration are in blue, and the final bundle-adjustment results are shown in red. Our method establishes reliable correspondences and accurately estimates camera poses for 3D reconstruction. For quantitative evaluation, please refer to Table 6
@@ D.2 Network architecture of SteerPose
Fig. 7 illustrates the network architecture of the proposed SteerPose. The input consists of 2D positions of $N$ joints in the image and three parameters representing the rotation by the Rodrigues vector. Each of the $N$ 2D joint positions and the rotation vector are projected onto the 32-dimensional vectors by MLP layers.
Each of the 32-dimensional vectors is combined with a positional encoding by adding them element-wise. Here we used a look-up embedding with $N+1$ tokens for positional encoding. The embeddings with the positional encoding are used as input tokens of the transformer encoder with five layers. The output tokens of the transformer encoder are averaged together to form a single 32-dimensional vector by mean pooling and then transformed to $N$ 2D joint positions.

### v2

@@ 3.2 Joint Extrinsic Camera Calibration and Matching
The geometric consistency loss $\mathcal{L}_{\mathrm{geom}}$ evaluates whether the combination of the current $R$ and the estimated matches yields a valid solution of the relative translation, after enforcing collinearity and coplanarity constraints  [35, 19]. For all NN corresponding 2D pose pairs under current rotation $R$, we transform their positions into a normalized coordinate system. By stacking linear equations derived from collinearity and coplanarity constraints [19], we can obtain a linear system of the form $A\begin{bmatrix}x_{1}&\dots&x_{N}&\bm{t}\end{bmatrix}^{\top}=\bm{0}$, where unknowns $x_{i}$ are 3D poses recovered from the collinearity constraint and $\bm{t}$ is determined up to scale and sign ambiguity. This sign ambiguity is resolved via a chirality check [35].
@@ 5.1 Experimental Setups
[…]
@@
[…]
@@
[…]
@@ 5.2 Comparison with Existing Methods
Figure 4: Visualization for multi-view calibration results. The ground-truth cameras are in black, the initial camera poses after registration are in blue, and the final bundle-adjustment results are shown in red. Our method establishes reliable correspondences and accurately estimates camera poses for 3D reconstruction. For quantitative evaluation, please refer to Table 6.
@@ D.2 Network architecture of SteerPose
Fig. 7 illustrates the network architecture of the proposed SteerPose. The input consists of 2D positions of NN joints in the image and three parameters representing the rotation by the Rodrigues vector. Each of the NN 2D joint positions and the rotation vector are projected onto the 32-dimensional vectors by MLP layers.
Each of the 32-dimensional vectors is combined with a positional encoding by adding them element-wise. Here we used a look-up embedding with $N+1$ tokens for positional encoding. The embeddings with the positional encoding are used as input tokens of the transformer encoder with five layers. The output tokens of the transformer encoder are averaged together to form a single 32-dimensional vector by mean pooling and then transformed to NN 2D joint positions.